2(-3a=5)=4(a+4)

ii know the answer is 12 but idk how to get it

Answer by mathman409(67)

YOU CAN'T HAVE TWO EQUAL SIGNS!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
You might have meant a plus sign. so, i'll solve it that way.
2(-3a+5)=4(a+4)
Step#1 OPEN THE PARENTHESES(do the multiplication to every number)
2(-3a+5)=4(a+4)
-6a+10=4a+16
Step#2 Write in standard form(put all variables in one side of the equal sign and the other numbers in the other)
-6a+10=4a+16
10=4a+16+6a
10-16=4a+6a
Step#3 SOLVE
10-16=4a+6a
-6=10a
-.6=a